OpenStudy (anonymous):

if you make 10 hr how much would you make a day

OpenStudy (anonymous):

What's the income?

OpenStudy (anonymous):

Is it 10 per hour?

OpenStudy (anonymous):

@King_Endy your question makes no sense. Is there more to it?

OpenStudy (anonymous):

My bad i meant @cryaotic

OpenStudy (anonymous):

sorry typed it wrong

OpenStudy (jdoe0001):

10hrs? heheh, well, then you just worked 10 hours =)

OpenStudy (anonymous):

On a side note, Cry all the way... and its fine.

OpenStudy (anonymous):

Can you retype it then?

OpenStudy (jdoe0001):

on the other hand... I dunno how to make hours... I do notice them pass by...

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

not enough to make your parents proud

OpenStudy (anonymous):

if you make 10hr how much do you make in a day

OpenStudy (jdoe0001):

10hrs? we're back to square one :|

OpenStudy (anonymous):

10/24 = ?

OpenStudy (jdoe0001):

if you work 10hrs, well then youb just worked 10hrs, hopefully for a good payrate

OpenStudy (anonymous):

Do you not know how to divide?

OpenStudy (anonymous):

@cryaotic Either your teacher made a stupid question or you dont know hoe to type your question. But stupid questions will get stupid answers

OpenStudy (anonymous):

He (or she) is asking for the amount of the day they're using.... 10/ 24, then?

OpenStudy (jdoe0001):

@cryaotic we get the hours... we just don't have a payrate to use say if you work 10hrs and you make $5 per hour, then you made in 1 day 5*10 if you work 10hrs and you get paid ... $20 per hour, then you made 20* 10 and so on

OpenStudy (anonymous):

Or am I off?

OpenStudy (anonymous):

I feel like @cryaotic is a guy. dont know why i just do. sorry of topic

OpenStudy (anonymous):

i make 15 dollars an hour

OpenStudy (jdoe0001):

$15... not so bad

OpenStudy (anonymous):

15/ 10 = x/24 ?

OpenStudy (jdoe0001):

@cryaotic so, you're after 10hrs of work for $15 per hour?

OpenStudy (anonymous):

15*10=?

OpenStudy (anonymous):

But the question asked for how much he made in a day @vicleon6

OpenStudy (anonymous):

So out of 24 hrs, how much did he make?

OpenStudy (anonymous):

so would i do 10*24 then /10?

OpenStudy (jdoe0001):

well, the day has 24hrs, doesn't mean he worked from midnight to midnight

OpenStudy (anonymous):

*15

OpenStudy (anonymous):

Then were did the 10 hours come from.

OpenStudy (anonymous):

In all, if its 15 in a total of 10 (15/10), how much will they make in 24 (x/24)?

OpenStudy (anonymous):

what exactly is the question?

OpenStudy (jdoe0001):

@cryaotic we don't even know the specific question to begin with

OpenStudy (anonymous):

Exactly!!! @jdoe0001

OpenStudy (anonymous):

From the 10 hours needed to make the 15. And true, @cryaotic

OpenStudy (anonymous):

i work for 10hrs i make 15$ an hr how much do i make a day

OpenStudy (anonymous):

Ya so its just 10*15 isn't it?

OpenStudy (anonymous):

But if its 10*15, how much would it be out of the whole day?

OpenStudy (anonymous):

Sorry for arguing, but it feels like an o

OpenStudy (anonymous):

"out of what" problem

OpenStudy (anonymous):

Well the question is saying he worked a 10 hr day and gets paid $15 an hr. so you would just multiply your two numbers.

OpenStudy (anonymous):

At least that's how I learned it.

OpenStudy (anonymous):

Ah. That could work. Sorry to argue. :(